# The General Position Problem in Graph Theory: A Survey

### v3

@@ 3.7 Steiner general position
[…]
@@
[64, Theorem 3.2] If $n\geq 3$ and $k\in\{2,\ldots,n-1\}$, then
@@
In [64, Theorem 4.2], a formula for $\operatorname{sgp}_{k}$ of the join $G\vee H$ of graphs $G$ and $H$ is given in terms of two related invariants which we do not explain here. Instead, we just mention that the formula can be simplified in many cases. We extract the following two formulas from a list of five formulas given in [64, Corollary 4.4]:
@@
The paper [64] further introduces bounds and exact results for the lexicographic product of graphs, for split graphs, and proves in [64, Theorem 6.2] that $\operatorname{sgp}_{k}(P_{\infty}\operatorname{\,\square\,}P_{\infty})\geq 2k$.
@@ 3.8 Detour position problem
The authors claim in [32, Theorem 3] that any longest path in a graph can contain at most two vertices of a detour position set, which would imply the upper bound $n-D+1$, where $D$ is the length of a longest path in $D$ (known as the detour diameter). However, the proof assumes that a subpath of a detour is also a detour, which is not true in general, so at present we regard the $n-D+1$ bound as a conjecture.
For any tree $T$, the longest paths and shortest paths coincide, so naturally $\operatorname{dp}(T)=\operatorname{gp}(T)=\ell(T)$. The paper discusses the relation between the detour position number and the detour hull number and also shows that for any $k\geq 2$ there is a graph with $\operatorname{dp}(G)=k$ and any feasible combination of detour radius and detour diameter. It is shown that a graph has detour position number $n-1$ if and only if it is $K_{3}$ or a star, and detour position number $n-2$ if and only if it is a star plus an edge or a double star (although both of these results depend on the $n-D+1$ bound).
@@ 4.1 Monophonic position sets
[113, Lemma 3.1] Let $G$ be a connected graph and $M\subseteq V(G)$ be a monophonic position set of $G$. Then $G[M]$ is a disjoint union of $k\geq 1$ cliques $G[M]=\bigcup_{i=1}^{k}W_{i}$. If $k\geq 2$, then for $i\in[k]$ any two vertices of $W_{i}$ have a common neighbour in $G\setminus M$.
@@
[113, Conjecture 2.9] The largest possible monophonic position number of a cubic graph with order $n$ is $\frac{n}{3}+O(n)$.
@@
[…]
@@
[…]
@@
[122, Conjecture 4.5] For any $n\geq 11$, there is a circulant graph with order $n$, monophonic position number two and diameter two.
The monophonic position numbers of the Cartesian and lexicographic product graphs are discussed in [21]. For the Cartesian product, as noted in [21, Observation 3.1, Corollary 3.6], we have that for any graphs $G$ and $H$,
@@
As these bounds coincide for products of paths and cycles, we have $\operatorname{mp}(P_{m}\operatorname{\,\square\,}P_{n})=\operatorname{mp}(P_{n}\operatorname{\,\square\,}C_{r})=\operatorname{mp}(C_{r}\operatorname{\,\square\,}C_{s})=2$, see [21, Corollary 3.7]. As reported in [21, Corollary 3.8], if $H$ is a connected graph and $n\geq\operatorname{mp}(H)$, then $\operatorname{mp}(K_{n}\operatorname{\,\square\,}H)=n.$ This in particular yields $\operatorname{mp}(K_{n}\operatorname{\,\square\,}P_{m})=n$ for $n\geq 2$, $\operatorname{mp}(K_{n}\operatorname{\,\square\,}C_{m})=n$ for $n\geq 3$, and $\operatorname{mp}(K_{n}\operatorname{\,\square\,}K_{m})=\max\{n,m\}$.
@@
[21, Theorem 3.16] If $G$ and $H$ are connected graphs, then
@@
[21, Theorem 3.18] further provides improved bounds for the monophonic position number of Cartesian products of triangle-free graphs.
@@
[21, Theorem 3.18] If $G$ and $H$ are connected triangle-free graphs of order at least three, then
@@
For a monophonic position set $M$ of $G$ we denote the components of $G[M]$ by $A_{1},A_{2},\ldots,A_{k},B_{1},\dots,B_{r}$, where $|A_{i}|\geq 2$ for each $i\in[k]$ and $|B_{j}|=1$ for each $j\in[r]$. Also we fix $n_{M}=\sum_{i=1}^{k}|A_{i}|$ and write $r_{M}=r$ to emphasise that $r$ is a function of $M$. Then for any monophonic position set $M$ of $G$ we have $|M|=n_{M}+r_{M}$. Using this notation, [21, Theorem 4.4] provides a formula for the monophonic position number of the lexicographic product of arbitrary graphs. The result reads as follows.
@@
[21, Theorem 4.4] Let $G$ be a connected graph of order at least two and let $\mathcal{M}$ be the collection of all monophonic position sets of $G$. Then
@@
[113, Theorem 6.2] The Monophonic position set problem is NP-hard.
@@ 4.2 All paths position number
[…]
@@ 5.1 Mobile position sets
As any pair of robots is trivially in general position, we have $\operatorname{Mob_{gp}}(G)\geq 2$ for any graph with $n(G)\geq 2$ (however, there is currently no characterisation of graphs with mobile general position number two). By way of illustration, consider $K_{n_{1},\dots,n_{t}}$, $n_{1}\geq 2$, $t\geq 2$. If at any point there are two or more robots in a partite set $V_{i}$, then there can be no robot in a set $V_{j}$, where $i\neq j$. Hence in this case for there to be an available legal move there can be only two robots on the graph. Thus if there are at least three robots on the graph, then they must all lie in different partite sets, so that there are at most $t$ robots. If there are $t$ robots, then any move would result in two robots in the same partite set, whereas it is easily seen that $t-1$ robots can visit all vertices of the graph whilst remaining in different partite sets. Thus the mobile general position number of this graph is $\max\{2,t-1\}$ and it follows immediately that for any $2\leq a\leq b$ there is a graph $G$ with $\operatorname{Mob_{gp}}(G)=a$ and $\operatorname{gp}(G)=b$.
@@
The paper [63] begins by discussing separable graphs, i.e. graphs containing a cut-vertex. Let $G$ be a graph with cut-vertex $v$. Denote the components of $G-v$ by $H_{1},\dots,H_{r}$ and for $i\in[r]$ the subgraph induced by $V(H_{i})\cup\{v\}$ by $G_{i}$. At some point a robot must visit the cut-vertex $v$, so at this point all robots must be contained within some $G_{i}$. Hence at any point at most two components of $G-v$ can contain robots, and if there are robots in two such components, then one of the components contains just one robot; it follows that there is a component $H_{j}$ that contains all the robots with one exception, a robot $R$ that traverses the rest of the graph. This is the source of an interesting mistake in Lemma 2.1.ii) and Corollary 2.2 of [63], which was pointed out by Ethan Shallcross. It was assumed that whilst $R$ is visiting other components, the robots remaining in $H_{j}$ must form a mobile general position set, which would imply that $\operatorname{Mob_{gp}}(G)=\operatorname{Mob_{gp}}(G_{k})$ for some $k\in[r]$. In fact, this is not the case, as can be seen by considering a cycle $C_{4}$ with a leaf attached to two adjacent vertices of the cycle. It remains an open question how much larger $\operatorname{Mob_{gp}}(G)$ can be than $\max\{\operatorname{Mob_{gp}}(G_{i}):\ i\in[r]\}$.
@@
[107, Theorem 3.1] The mobile general position number of any non-complete graph $G$ with clique number $\omega(G)\geq 3$ is bounded below by
@@
[62, Proposition 2.1] For any connected graphs $G$ and $H$ of order at least two, the following hold.
@@
Both lower bounds in Proposition 5.4 are sharp, as demonstrated by the following example given in [62, Proposition 2.2]: if $r,s\geq 2$, then
@@
The paper [62] gives exact values of $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{m})$ $(n,m\geq 1)$, $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})$ $(n,m\geq 3)$, $\operatorname{Mob_{gp}}(C_{n}\operatorname{\,\square\,}K_{2})$ $(n\geq 3)$, $\operatorname{Mob_{gp}}(C_{4}\operatorname{\,\square\,}P_{s})=3$ $(r\geq 3)$, $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})$, and $(r=9\ {\rm or}\ r\geq 11\ {\rm and}\ s\geq 5)$, and proves that $\operatorname{Mob_{gp}}(T\operatorname{\,\square\,}K_{2})=\ell(T)$ for any tree $T$ with order at least three [62, Theorem 3.1]. For the corona product, lower and upper bounds are given on $\operatorname{Mob_{gp}}(G\odot H)$ in [62, Theorem 4.1] and it is proved that $\operatorname{Mob_{gp}}(C_{n}\odot K_{1})=\left\lceil\frac{n}{2}\right\rceil+1$ [62, Theorem 4.3]. Bounds on the mobile general position number of joins $G\vee H$ are also given.
@@ 5.2 Lower position numbers
[…]
Whilst Dudeney’s puzzle involves any straight line in the plane, Gardner points out that this ‘minimum’ puzzle remains difficult even if we restrict attention to the set of lines corresponding to rows, columns and diagonals; this is often called the Queens version of the problem. This restricted version was treated in [30], in which the authors show using Combinatorial Nullstellensatz that the answer is at least $n$, except if $n\equiv 3\pmod{4}$, in which case the answer is at least $n-1$. The authors also include some details of Gardner’s correspondence on this problem (which show that this result was proven earlier by John Harris, although his argument is not extant). It is shown in [89] that $n+1$ is a lower bound when $n\equiv 1\pmod{4}$.
@@
[101, Theorem 4.1] Let $G$ and $H$ be non-trivial, connected graphs. Then $\operatorname{gp}^{-}(G\operatorname{\,\square\,}H)=2$ if and only if one of the following conditions holds.
@@
[…]
@@
The article [37] also considers the largest possible size of a graph with order $n$ and given lower $\operatorname{gp}$-number $k$; in contrast to the extremal question for the ordinary general position number, this problem is completely solved  [37, Theorem 10], and the unique extremal graph is formed by deleting the edges of a copy of the star $S_{k}$ from a clique $K_{n}$. Regarding complexity, [37] shows that the independent domination problem can be polynomially reduced to the lower general position problem, so that the latter is NP-complete.
@@
[37, Theorem 4.2] The Lower General Position problem is NP-complete.
@@
[37, Theorem 6.1] For $a,b\geq 1$, there exists a graph $G$ with $\operatorname{mp}^{-}(G)=a$ and $\operatorname{gp}^{-}(G)=b$ if and only if $a=b$, $2\leq a<b$ or $3\leq b<a$.
@@
[77, Theorem 2.6] If $k\geq 2$ and $n_{1}\geq 2,\ldots,n_{k}\geq 2$, then
@@
[77, Theorem 4.6] If $G=K_{m_{1},\dots,m_{r}}$ and $H=K_{n_{1},\dots,n_{s}}$, where $r,s\geq 2$, $m_{1}\geq\dots\geq m_{r}$, $n_{1}\geq\dots\geq n_{s}$, and $m_{1},n_{1}\geq 2$, then
@@
[77, Conjecture 3.3] Every graph has a terminal set.

### v4

@@ 3.7 Steiner general position
Let $G$ be a connected graph and $X\subseteq V(G)$. The Steiner distance $d_{G}(X)$ of $X$ is the minimum size of a connected subgraph of $G$ containing $X$ [27]. Such a subgraph is clearly a tree and is called a Steiner $X$-tree. For a positive integer $k$, we say that $A\subseteq V(G)$ is a $k$-Steiner general position set if for every $B\subseteq A$, $|B|=k$, and for every Steiner $B$-tree $T_{B}$, it follows that $V(T_{B})\cap A=B$. That is, $A$ is a $k$-Steiner general position set if no $k+1$ distinct vertices from $A$ lie on a common Steiner $B$-tree, where $B\subseteq A$ and $|B|=k$. Clearly, if $|A|\leq k$, then $A$ is $k$-Steiner general position set. The $k$-Steiner general position number $\operatorname{sgp}_{k}(G)$ of $G$ is the cardinality of a largest $k$-Steiner general position set in $G$. Note that $\operatorname{sgp}_{2}(G)=\operatorname{gp}(G)$.
In [73, Proposition 2.2] it was observed that if $G$ is a graph and $k\in\{2,\ldots,n(G)-1\}$, then $\operatorname{sgp}_{k}(G)=n(G)$ if and only if $G$ is (n(G)k+1)(n(G)-k+1)-connected. If $T$ is a tree, then $\operatorname{sgp}_{k}(T)=\ell(T)$ when $k\leq\ell(T)$, and $\operatorname{sgp}_{k}(T)=k$ when $k>\ell(T)$. The corresponding result for cycles reads as follows.
@@
[73, Theorem 3.2] If $n\geq 3$ and $k\in\{2,\ldots,n-1\}$, then
@@
In [73, Theorem 4.2], a formula for $\operatorname{sgp}_{k}$ of the join $G\vee H$ of graphs $G$ and $H$ is given in terms of two related invariants which we do not explain here. Instead, we just mention that the formula can be simplified in many cases. We extract the following two formulas from a list of five formulas given in [73, Corollary 4.4]:
@@
The paper [73] further introduces bounds and exact results for the lexicographic product of graphs, for split graphs, and proves in [73, Theorem 6.2] that $\operatorname{sgp}_{k}(P_{\infty}\operatorname{\,\square\,}P_{\infty})\geq 2k$.
@@ 3.8 Detour position problem
The authors claim in [36, Theorem 3] that any longest path in a graph can contain at most two vertices of a detour position set, which would imply the upper bound nD+1n-D+1, where $D$ is the length of a longest path in $D$ (known as the detour diameter). However, the proof assumes that a subpath of a detour is also a detour, which is not true in general, so at present we regard the nD+1n-D+1 bound as a conjecture.
For any tree $T$, the longest paths and shortest paths coincide, so naturally $\operatorname{dp}(T)=\operatorname{gp}(T)=\ell(T)$. The paper discusses the relation between the detour position number and the detour hull number and also shows that for any $k\geq 2$ there is a graph with $\operatorname{dp}(G)=k$ and any feasible combination of detour radius and detour diameter. It is shown that a graph has detour position number n1n-1 if and only if it is $K_{3}$ or a star, and detour position number n2n-2 if and only if it is a star plus an edge or a double star (although both of these results depend on the nD+1n-D+1 bound).
@@ 4.1 Monophonic position sets
[129, Lemma 3.1] Let $G$ be a connected graph and $M\subseteq V(G)$ be a monophonic position set of $G$. Then $G[M]$ is a disjoint union of $k\geq 1$ cliques $G[M]=\bigcup_{i=1}^{k}W_{i}$. If $k\geq 2$, then for $i\in[k]$ any two vertices of $W_{i}$ have a common neighbour in $G\setminus M$.
@@
[129, Conjecture 2.9] The largest possible monophonic position number of a cubic graph with order $n$ is $\frac{n}{3}+O(n)$.
@@
[129, Conjecture 3.4] For sufficiently large $g$, the monophonic position number of a $(d,g)$-cage $G$ satisfies $\operatorname{mp}(G)<d$.
@@
The final section of [140] classifies the possible diameters of a graph with $\operatorname{mp}$-number $a$. If $a\geq 3$, then all diameters in the range $2\leq D\leq n-a+1$ are possible and if $a=2$, then the possible diameters are D=n1D=n-1, the integers $D$ in the range $3\leq D\leq\lfloor\frac{n}{2}\rfloor$ and, if $n\in\{3,4,5,8\}$ or $n\geq 11$, diameter $D=2$. The case of graphs with $\operatorname{mp}$-number two and diameter two is rather interesting, and a new graph operation is defined that, given such a graph with order $n$, will produce a new graph with the same properties and order either $3n+2$, $3n+1$ or $3n$. An example of this operation applied to $C_{4}$ can be seen in Figure 6. This construction may be worthy of further study.
@@
[140, Conjecture 4.5] For any $n\geq 11$, there is a circulant graph with order $n$, monophonic position number two and diameter two.
The monophonic position numbers of the Cartesian and lexicographic product graphs are discussed in [24]. For the Cartesian product, as noted in [24, Observation 3.1, Corollary 3.6], we have that for any graphs $G$ and $H$,
@@
As these bounds coincide for products of paths and cycles, we have $\operatorname{mp}(P_{m}\operatorname{\,\square\,}P_{n})=\operatorname{mp}(P_{n}\operatorname{\,\square\,}C_{r})=\operatorname{mp}(C_{r}\operatorname{\,\square\,}C_{s})=2$, see [24, Corollary 3.7]. As reported in [24, Corollary 3.8], if $H$ is a connected graph and $n\geq\operatorname{mp}(H)$, then $\operatorname{mp}(K_{n}\operatorname{\,\square\,}H)=n.$ This in particular yields $\operatorname{mp}(K_{n}\operatorname{\,\square\,}P_{m})=n$ for $n\geq 2$, $\operatorname{mp}(K_{n}\operatorname{\,\square\,}C_{m})=n$ for $n\geq 3$, and $\operatorname{mp}(K_{n}\operatorname{\,\square\,}K_{m})=\max\{n,m\}$.
@@
[24, Theorem 3.16] If $G$ and $H$ are connected graphs, then
@@
[24, Theorem 3.18] further provides improved bounds for the monophonic position number of Cartesian products of triangle-free graphs.
@@
[24, Theorem 3.18] If $G$ and $H$ are connected triangle-free graphs of order at least three, then
@@
For a monophonic position set $M$ of $G$ we denote the components of $G[M]$ by $A_{1},A_{2},\ldots,A_{k},B_{1},\dots,B_{r}$, where $|A_{i}|\geq 2$ for each $i\in[k]$ and $|B_{j}|=1$ for each $j\in[r]$. Also we fix $n_{M}=\sum_{i=1}^{k}|A_{i}|$ and write $r_{M}=r$ to emphasise that $r$ is a function of $M$. Then for any monophonic position set $M$ of $G$ we have $|M|=n_{M}+r_{M}$. Using this notation, [24, Theorem 4.4] provides a formula for the monophonic position number of the lexicographic product of arbitrary graphs. The result reads as follows.
@@
[24, Theorem 4.4] Let $G$ be a connected graph of order at least two and let $\mathcal{M}$ be the collection of all monophonic position sets of $G$. Then
@@
[129, Theorem 6.2] The Monophonic position set problem is NP-hard.
@@ 4.2 All paths position number
It turns out that the all-path position problem is intimately connected to connectivity. For a 2-connected graph with order $n(G)\geq 2$ the all-path position number is two (this can be shown using the Fan Lemma), whereas for a graph $G$ with a cut-vertex the all-path position number is equal to the number of leaf blocks of $G$ (i.e. the number of blocks of $G$ containing a single cut-vertex of $G$). This result was also proved independently in unpublished work by Sumaiyah Boshar [15]. We suggest that it would be of interest to consider the more general no-$k$-in-line problem for all paths.
@@ 5.1 Mobile position sets
As any pair of robots is trivially in general position, we have $\operatorname{Mob_{gp}}(G)\geq 2$ for any graph with $n(G)\geq 2$ (however, there is currently no characterisation of graphs with mobile general position number two). By way of illustration, consider $K_{n_{1},\dots,n_{t}}$, $n_{1}\geq 2$, $t\geq 2$. If at any point there are two or more robots in a partite set $V_{i}$, then there can be no robot in a set $V_{j}$, where $i\neq j$. Hence in this case for there to be an available legal move there can be only two robots on the graph. Thus if there are at least three robots on the graph, then they must all lie in different partite sets, so that there are at most $t$ robots. If there are $t$ robots, then any move would result in two robots in the same partite set, whereas it is easily seen that t1t-1 robots can visit all vertices of the graph whilst remaining in different partite sets. Thus the mobile general position number of this graph is $\max\{2,t-1\}$ and it follows immediately that for any $2\leq a\leq b$ there is a graph $G$ with $\operatorname{Mob_{gp}}(G)=a$ and $\operatorname{gp}(G)=b$.
@@
The paper [72] begins by discussing separable graphs, i.e. graphs containing a cut-vertex. Let $G$ be a graph with cut-vertex $v$. Denote the components of GvG-v by $H_{1},\dots,H_{r}$ and for $i\in[r]$ the subgraph induced by $V(H_{i})\cup\{v\}$ by $G_{i}$. At some point a robot must visit the cut-vertex $v$, so at this point all robots must be contained within some $G_{i}$. Hence at any point at most two components of GvG-v can contain robots, and if there are robots in two such components, then one of the components contains just one robot; it follows that there is a component $H_{j}$ that contains all the robots with one exception, a robot $R$ that traverses the rest of the graph. This is the source of an interesting mistake in Lemma 2.1.ii) and Corollary 2.2 of [72], which was pointed out by Ethan Shallcross. It was assumed that whilst $R$ is visiting other components, the robots remaining in $H_{j}$ must form a mobile general position set, which would imply that $\operatorname{Mob_{gp}}(G)=\operatorname{Mob_{gp}}(G_{k})$ for some $k\in[r]$. In fact, this is not the case, as can be seen by considering a cycle $C_{4}$ with a leaf attached to two adjacent vertices of the cycle. It remains an open question how much larger $\operatorname{Mob_{gp}}(G)$ can be than $\max\{\operatorname{Mob_{gp}}(G_{i}):\ i\in[r]\}$.
@@
[121, Theorem 3.1] The mobile general position number of any non-complete graph $G$ with clique number $\omega(G)\geq 3$ is bounded below by
@@
[71, Proposition 2.1] For any connected graphs $G$ and $H$ of order at least two, the following hold.
@@
Both lower bounds in Proposition 5.4 are sharp, as demonstrated by the following example given in [71, Proposition 2.2]: if $r,s\geq 2$, then
@@
The paper [71] gives exact values of $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{m})$ $(n,m\geq 1)$, $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})$ $(n,m\geq 3)$, $\operatorname{Mob_{gp}}(C_{n}\operatorname{\,\square\,}K_{2})$ $(n\geq 3)$, $\operatorname{Mob_{gp}}(C_{4}\operatorname{\,\square\,}P_{s})=3$ $(r\geq 3)$, $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})$, and $(r=9\ {\rm or}\ r\geq 11\ {\rm and}\ s\geq 5)$, and proves that $\operatorname{Mob_{gp}}(T\operatorname{\,\square\,}K_{2})=\ell(T)$ for any tree $T$ with order at least three [71, Theorem 3.1]. For the corona product, lower and upper bounds are given on $\operatorname{Mob_{gp}}(G\odot H)$ in [71, Theorem 4.1] and it is proved that $\operatorname{Mob_{gp}}(C_{n}\odot K_{1})=\left\lceil\frac{n}{2}\right\rceil+1$ [71, Theorem 4.3]. Bounds on the mobile general position number of joins $G\vee H$ are also given.
@@ 5.2 Lower position numbers
In 1976 the famous recreational mathematician (and modern day Henry Dudeney) Martin Gardner posed the following ‘worst-case’ version of Dudeney’s chessboard puzzle in his column in Scientific American [54]: what is the smallest number of pawns that can be placed on an $n\times n$ chessboard, such that adding any new pawn creates three-in-a-line? The results can be found as sequence A277433 in the Online Encyclopedia of Integer Sequences [101]. Note that the pawns in this set do not necessarily themselves have the no-three-in-line property; when the no-three-in-line property is required, this is equivalent to a smallest maximal no-three-in-line set (see sequence A219760 in OEIS). This problem is also mentioned in [2]. The problem is taken up, under the name geometric dominating sets, in the article [4], which presents a lower bound $\Omega(n^{2/3})$ and an upper bound $2\left\lceil\frac{n}{2}\right\rceil$ (and the discrete torus is also investigated). When the pawns are required to be no-three-in-line (which the authors call independent geometric dominating sets), they find the solutions for $n\leq 12$, but do not improve on the trivial $2n$ upper bound.
Whilst Dudeney’s puzzle involves any straight line in the plane, Gardner points out that this ‘minimum’ puzzle remains difficult even if we restrict attention to the set of lines corresponding to rows, columns and diagonals; this is often called the Queens version of the problem. This restricted version was treated in [34], in which the authors show using Combinatorial Nullstellensatz that the answer is at least $n$, except if $n\equiv 3\pmod{4}$, in which case the answer is at least n1n-1. The authors also include some details of Gardner’s correspondence on this problem (which show that this result was proven earlier by John Harris, although his argument is not extant). It is shown in [100] that $n+1$ is a lower bound when $n\equiv 1\pmod{4}$.
@@
[115, Theorem 4.1] Let $G$ and $H$ be non-trivial, connected graphs. Then $\operatorname{gp}^{-}(G\operatorname{\,\square\,}H)=2$ if and only if one of the following conditions holds.
@@
For $n\geq 12$, a lower $\operatorname{gp}$-set is the set of six 22-subsets of four symbols from $[n]$.
@@
The article [41] also considers the largest possible size of a graph with order $n$ and given lower $\operatorname{gp}$-number $k$; in contrast to the extremal question for the ordinary general position number, this problem is completely solved  [41, Theorem 10], and the unique extremal graph is formed by deleting the edges of a copy of the star $S_{k}$ from a clique $K_{n}$. Regarding complexity, [41] shows that the independent domination problem can be polynomially reduced to the lower general position problem, so that the latter is NP-complete.
@@
[41, Theorem 4.2] The Lower General Position problem is NP-complete.
@@
[41, Theorem 6.1] For $a,b\geq 1$, there exists a graph $G$ with $\operatorname{mp}^{-}(G)=a$ and $\operatorname{gp}^{-}(G)=b$ if and only if $a=b$, $2\leq a<b$ or $3\leq b<a$.
@@
[88, Theorem 2.6] If $k\geq 2$ and $n_{1}\geq 2,\ldots,n_{k}\geq 2$, then
@@
[88, Theorem 4.6] If $G=K_{m_{1},\dots,m_{r}}$ and $H=K_{n_{1},\dots,n_{s}}$, where $r,s\geq 2$, $m_{1}\geq\dots\geq m_{r}$, $n_{1}\geq\dots\geq n_{s}$, and $m_{1},n_{1}\geq 2$, then
@@
[88, Conjecture 3.3] Every graph has a terminal set.